OpenStudy (anonymous):

Solve the radical equation, and check all proposed solutions:

OpenStudy (anonymous):

\[\sqrt{7x+44}=x\]

hero (hero):

Square both sides first. Let me know what you get afterwards.

OpenStudy (anonymous):

\[7x+44=x^{2}\]

hero (hero):

Subtract 7x from both sides. Then subtract 44 from both sides. Let me know what you get afterwards.

OpenStudy (anonymous):

\[0=x ^{2}-7x-44\]

hero (hero):

Split the middle term so that -7x = -11x + 4x

hero (hero):

\[0 = x^2 - 11x + 4x - 44\]

hero (hero):

Now factor by grouping

OpenStudy (anonymous):

(x-4)(x-11)?

hero (hero):

(x-4)(x-11) = 0 But you still have to solve for x

OpenStudy (anonymous):

how do I do that?

OpenStudy (anonymous):

Oh wait, it's (x+4)

hero (hero):

Yes, you're right

hero (hero):

x + 4 = 0 x - 11 = 0 Solve for x for both Only one x is the correct solution

OpenStudy (anonymous):

How would you find x for things like (5x+3)(5x+9)=0?

hero (hero):

Well, actually, both solutions work now that x = -4

hero (hero):

Again do the same thing 5x + 3 = 0 5x + 9 = 0 Solve for x

hero (hero):

Basically, you're applying what is called zero product property.

hero (hero):

`Zero Product Property If the product of two numbers, a and b are zero, then a or b must equal zero:` In other words, at least one of those numbers will be equal to zero.

OpenStudy (anonymous):

No, I got 4 not -4 so it didn't work

OpenStudy (anonymous):

What would be the domain of this equation?